# MiM: Mask in Mask Self-Supervised Pre-Training for 3D Medical Image Analysis

### v1

@@ 3.3 Cross-level alignment
Fig. 4: Illustration of cross-level alignment module. We set level $L$ to 2 for better illustration. The Level-2 context representation and the corresponding patches of Level-1 volumes are regarded as positive pairs. The rest patches of the Level-1 volume are considered negative pairs.
@@
We aim to maximize the cosine similarity between the context vector $c_{i}$ and the positive patches $p$ and minimize the cosine similarity between the context vector $c_{i}$ and the negative patches $p_{j}$ from the rest patches of the volumes. Thus, we apply the contrastive loss [14] to implement this goal as follows,
@@ 3.4 Backbone
Previous hybrid transformers, such as nnFormer [88] and Swin-UNETR [28, 62], aim to produce pyramid features. However, they suffer from complexity, needing to process all visual tokens for Masked Image Modeling, which is inefficient [22] and hinders scalability. We extend the backbone of [22] from 2D to 3D medical images by adding an extra dimension channel i.e., depth. This modification allows us to process visual tokens only on transformer layers, enhancing computational efficiency and scalability. The ablation study confirming the effectiveness of this approach is presented in Section 4.7.1.
Overall objective function. Our proposed MiM framework, as illustrated in Algorithm 1, introduces a hierarchically designed approach to 3D medical image representation learning through multi-level reconstruction $\mathcal{L}_{R}$ and cross-level alignment $\mathcal{L}_{C}$. We empirically set the level of masked volume to $L=3$ in this paper since the three-level masked volume can provide a good balance between the representation learning and computational efficiency. Abaltion study of the level of masked volume $L$ is on the Section 4.7.1. Specifically, the multi-level reconstruction loss in Eq. 2 can be further expanded as follows:
@@
Then, since cross-level alignment loss applied between adjacent level volumes, Eq. 6 can be also further expanded as follows:
@@
where the hyperparameters $\alpha$ are used to balance the relative contributions of these two kinds of losses. We empirically set the $\alpha$ to 0.1 based on our experiment results. The ablation study of hyperparameter $\alpha$ presents in Section 4.7.2.
input :
1 $\mathcal{D}$, $\tau$     dataset and the transformations;
2 $\theta$, $s_{\theta}$    initial parameters for encoder, encoder;
3 $\xi$, $d_{\xi}$    initial parameters for decoder, decoder;
4 $\iota$, $\mathcal{H}$    initial parameters for aggregation and projector heads;
5 $\operatorname{optimizer}$    optimizer, updated parameters via gradient backwards;
6 $K$, $B$, $\mu$    pre-training steps, batch size and learning rate; for $k\leftarrow 1$ to $K$ do
       // sample one batches from dataset
7       $\mathcal{B}$leftarrow $\left\{\left\{x\right\}\sim\mathcal{D}\right\}_{i=1}^{B}$ for i, $x\in\mathcal{B}$ do
8
9            Apply multi-level masked volume generation for $x$ to obtain $x^{1}$, $x^{2}$ and $x^{3}$;
10             Apply Eq.(2) for $x^{1}$ to obtain $\mathcal{L}_{R}^{1}$ with $s_{\theta}$ and $d_{\xi}$;
11             Apply Eq.(2) for $x^{2}$ to obtain $\mathcal{L}_{R}^{2}$ with $s_{\theta}$ and $d_{\xi}$;
12             Apply Eq.(2) for $x^{3}$ to obtain $\mathcal{L}_{R}^{3}$ with $s_{\theta}$ and $d_{\xi}$;
13             Apply Eq.(3) to obtain $\mathcal{L}^{R}$ with $s_{\theta}$ and $d_{\xi}$;
14             Apply Eq.(6) for $x^{2}$ and $x^{1}$ to obtain $\mathcal{L}_{C}^{1,2}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
15             Apply Eq.(6) for $x^{3}$ and $x^{2}$ to obtain $\mathcal{L}_{C}^{2,3}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
16             Apply Eq.(7) to obtain $\mathcal{L}^{C}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
17             Apply Eq.(10) to obtain the final loss $\mathcal{L}$
18      $\delta\theta\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\theta}\mathcal{L}\right);$ $\delta\xi\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\xi}\mathcal{L}\right);$
19       $\delta\iota\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\iota}\mathcal{L}\right);$
20       $\theta\leftarrow\operatorname{optimizer}(\theta,\delta_{\theta},\mu);$ $\xi\leftarrow\operatorname{optimizer}(\xi,\delta_{\xi},\mu);$
21       $\iota\leftarrow\operatorname{optimizer}(\iota,\delta_{\iota},\mu);$
22
output : $s_{\theta}$ // the pre-trained backbone encoder
23
Algorithm 1 MiM: Mask in Mask Self-Supervised Pre-training framework
@@ 4 Experiments
Table 1: The details of each clinical dataset in our pre-training datasets.
@@
BTCV [43] Abdomen 24 1. Change the volumetric orientation based on the given settings, e.g., RAS
@@
LUNA16 [27] Chest 843 2. Clip intensity value to [-1000, 1000] and normalize to the range of [0, 1].
@@
FLARE23 Abdomen 4,000 3. Pad each dimension with 0 to meet the minimum shape of input.
@@
HNSCC [25] Head/Neck 1,071 4. Randomly crop the sub-volume with the specific shape $H_{1}\times W_{1}\times D_{1}$
@@
This section will commence by introducing the pre-training datasets, downstream datasets, and evaluation metrics. Subsequently, we will elaborate on the implementation details of our MiM. Lastly, we will present the experimental results of MiM in comparison to other SOTA methods on various tasks and datasets, along with an analysis of our proposed approach.
4.1 Datasets and Evaluation
Pre-training datasets. To conduct a fair comparison with previous works [16, 62, 34, 70, 78, 91, 90, 89, 94], we also carried out pre-training experiments on two public datasets, i.e., BTCV [43] and TCIA Covid19 [17], and combined them to form a new dataset named 1k. Additionally, to explore the scaling ability of our proposed method compared to previous state-of-the-art methods [16, 62], we collected eight publicly accessible 3D medical image datasets consisting of 10,502 CT scans to establish our pre-training datasets, which we named 10k. It is important to note that 10k is only used for exploring the scaling ability of our proposed method, while we mainly focus on the 1k dataset for fair comparison and analysis of our proposed method. Table 1 provides a summary of the sources of each collected dataset. Our pre-training process solely utilizes volumetric images without involving any annotations from any of these datasets.
@@ 4.1 Datasets and Evaluation
Downstream datasets. To evaluate the effectiveness of our MiM, we conduct experiments on thirteen public datasets, i.e, BTCV [43], MM-WHS [37], Spleen [58], Flare22 [50], Amos22 [37], MSD Task03 [58], MSD Task06 [58], MSD Task07 [58], MSD Task08 [58], MSD Task10 [58], Covid-19-20 [56], BrasTS 21 [58], and CC-CCII [82]. These datasets include segmentation and classification tasks, with the first ten datasets used for organ segmentation, the eleventh dataset used for lesion segmentation, the twelfth dataset used for tumor segmentation, and the last dataset used for disease classification. Table 2 provides a summary of the details of each clinical dataset used in our downstream tasks. It’s noted that BTCV [43] is used in pre-training, the other datasets are unseen in pre-training. Furthermore, to evaluate the cross-modality generalization ability, we transferred the model pre-trained on CT scans to the MRI dataset BrasTS 21 [58]. We adopted consistent settings as previous work [16, 34, 62, 94].
Table 2: The details of each clinical datasets in our downstream datasets.
Datasets
Regions of
Interest
Modality #Samples Pre-processing
BTCV [43] Chest CT 30
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.1, 0.1 and 0.5
MM-WHS [95] Heart CT 20
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [0, 1700]
Spleen [58] Spleen CT 41
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-57, 175]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
Amos22 [95] Chest CT 301
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.5 and 0.5
Flare22 [95] Chest CT 50
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.2, 0.2 and 0.1
MSD Task03 [58] Liver CT 125
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-21,189]
3. Random flip, rotation and intensities shifting with probabilities of 0.2, 0.2 and 0.1
MSD Task06 [58] Lung CT 63
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-1000, 1000]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.3 and 0.1
MSD Task07 [58] Pancreas CT 281
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-87, 199]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
MSD Task08 [58]
Hepatic
Vessel
CT 303
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [0, 230]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
MSD Task10 [58] Colon CT 126
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-57, 175]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
Covid-19-20 [56]
Covid-19
lesion
CT 189
[…]
2. Crop the foregrounds with $192\times 192\times 32$ and normalize via HU range [-1000, 500]
BraTS 21 [58] Brain tumor MRI 484
[…]
2. Crop the foregrounds with $224\times 224\times 120$ and normalize via z-score across channels
CC-CCII [82]
Covid-19
diseases
CT 4178
Crop the foregrounds with $384\times 384\times 32$ and normalize via min-max normalization
Evaluation metrics. Following [16, 62, 94], we utilized the Dice Similarity Coefficient (DSC) [^1] https://docs.monai.io/en/latest/metrics.html#mean-dice, and Normalized Surface Distance (NSD)  [^2] https://docs.monai.io/en/latest/metrics.html#monai.metrics.SurfaceDiceMetric to assess the performance on segmentation tasks. We then utilized Accuracy (ACC) and the Area Under the Curve (AUC)  [^3] https://docs.monai.io/en/latest/metrics.html#area-under-the-roc-curve for accessing performance on disease classification.
4.2 Implementation Details and Evaluation metrics
Implementation details. In the pre-training process, we followed the settings of previous works [16, 62, 70] and provided the details of our MiM pre-training settings in Table 3. We used the backbone of  [22] as the encoder for efficient token processing. As previous SSL [4, 94], the prediction head and projection head are implemented with MLP layers for aligning the dimension. In the downstream process, we used Swin-UNETR [28] for fine-tuning on segmentation tasks and Swin-ViT for fine-tuning on classification tasks, as per the settings of previous works [62, 34, 70]. We initialized the encoder of the network with learned parameters in the pre-training process and fine-tuned the overall network. We used the AdamW [42] optimizer with a cosine learning rate scheduler for all experiments. For inference on these datasets, we applied a slicing window inference with overlapping to enable a fair comparison with previous works [62, 70]. All methods were pre-trained on four NVIDIA 3090Ti and then fine-tuned on a single NVIDIA 3090Ti. It is important to note that for evaluating the pure effectiveness of our proposed method, we did not use any foundation model or post-processing techniques [38, 47].
@@ 4.2 Implementation Details and Evaluation metrics
MiM Level-1 $H^{1}\times W^{1}\times D^{1}$ $384\times 384\times 192$
MiM Level-2 $H^{2}\times W^{2}\times D^{2}$ $96\times 96\times 96$
MiM Level-3 $H^{3}\times W^{3}\times D^{2}$ $16\times 16\times 16$
MiM Resize after crop $h\times w\times d$ $96\times 96\times 96$
MiM $\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}$ 1.0, 1.0, 1.0, 0.1, 0.1
@@
Comparison methods. In our experiments, we compared our proposed MiM method with both General SSL and Medical SSL methods. First, we compare the typical SSL methods MoCov3 [15] and MAE [16, 30], since they represent the two mainstream SSL paradigms, i.e., Contrastive learning and Pretext task. We also report the results of SimCLR [10] and PL [85], Jigsaw [8] according to  [16, 94]. We further evaluate the performance of SimMIM [79], HPM [70], localMIM [16] and MCMAE [22], since they are related to our advanced hybrid MAE. We also compare with Adam [35] due to related to our hierarchical design of multi-granularity. In addition, we compared with most existing state-of-the-art medical SSL methods in our experiments.
4.3 Experiments on downstream tasks
4.3.1 Comparison on the BTCV dataset
Outperform existing methods. We first conducted experiments on the BTCV [43], and the results are presented in Table 4. Among the comparision methods, SimCLR [14], MoCov3 [15], DINO [4], localMIM [67], HPM [67], MAE3D [16], GL-MAE [94] adopted UNETR [29] as networks architecture. Most other methods, including our MiM, used Swin-UNETR [28] as per the settings of previous works [62].
@@ 4.3.1 Comparison on the BTCV dataset
Method Pre-traning dataset Dice Score(%) Avg
@@
UNETR [29] - 93.02 94.13 94.12 66.99 70.87 96.11 77.27 89.22 82.10 70.16 76.65 65.32 59.21 79.82
Swin-UNETR [28] - 94.06 93.54 93.80 65.51 74.60 97.09 75.94 91.80 82.36 73.63 75.19 68.00 61.11 80.53
with General SSL
SimCLR [10] 1k 92.79 93.04 91.41 49.65 50.99 98.49 77.92 85.56 80.58 64.37 67.16 59.04 48.99 73.85
MoCov3 [15] 1k 91.96 92.85 92.42 68.25 72.77 94.91 78.82 88.21 81.59 71.15 75.76 66.48 58.81 79.54
Jigsaw [8] 1k 94.62 93.41 93.55 75.63 73.21 95.71 80.80 89.41 84.78 71.02 79.57 65.68 60.22 81.35
[…]
SimMIM [79] 1k 95.51 93.61 93.49 67.91 73.50 96.46 81.15 89.78 84.86 72.45 75.70 66.89 64.46 81.41
@@
with 3D Medical SSL
ROT [61] 1k 91.75 93.13 91.62 65.09 76.55 94.21 86.16 89.74 83.08 71.13 81.55 67.90 63.72 81.20
Rubik++ [92] 1k 96.21 90.41 89.33 75.22 72.64 97.44 79.25 89.65 83.76 74.74 78.35 67.14 61.97 81.38
ModelGen [91] 1k 91.99 93.52 91.81 65.11 76.14 95.98 86.88 89.29 83.59 71.79 81.62 67.97 63.18 81.45
@@
GL-MAE [94] 1k 94.54 94.39 94.37 73.19 74.93 96.51 83.49 89.74 83.11 70.80 75.71 69.39 63.12 82.33
[…]
MAE3D [16] 1k 95.81 94.38 94.48 69.96 76.85 96.69 80.44 90.33 84.33 73.65 80.11 68.65 64.44 82.40
MiM 1k 96.05 94.58 94.53 75.73 77.36 97.03 83.23 90.37 87.64 74.97 82.62 71.02 68.80 84.46
@@
Fig. 5: Visualization of segmentation results on BTCV validation dataset [43]. We compared MiM with Unimiss [78], PCRLv2 [89], GVSL [34], Swin-UNETR [62] and MAE3D [16].
Remark. As shown in Table 4, we observed that General SSL methods performed worse than Medical SSL methods. Specifically, SimCLR [14] and MoCov3 [15] achieved only 73.85% and 79.54%, respectively. This is because these methods rely on a large batch size and negative samples to avoid trivial constants, which is impractical for 3D medical images. Additionally, the negative relationship between different images used in SimCLR [14, 11] and MoCo [31, 12, 15] is not suitable for 3D medical images. Jigsaw [8] and DINO [4] also achieved limited improvements. Our proposed method, MiM, outperformed MAE-based methods such as MAE3D [16], GL-MAE [94], localMIM [67], HPM [67], and MCMAE [22] by a significant margin. We concluded that General SSL methods are not very suitable for 3D medical images, and it is crucial to consider the characteristics of 3D medical images when designing SSL methods.

### v2

@@ 3.3 Cross-level alignment
Figure 4: Illustration of cross-level alignment module. We set level $L$ to 2 for better illustration.
@@
We aim to maximize cosine similarity between the context vector $c_{i}$ and the positive patches $p$ and minimize cosine similarity between the context vector $c_{i}$ and the negative patches $p_{j}$ from the rest patches of the volumes. Thus, we apply contrastive loss [29] to implement this goal as follows,
@@ 3.4 Backbone
While previous hybrid transformers like Swin Transformer [21] generate pyramid features, they process all tokens during Masked Image Modeling pre-training, leading to computational inefficiency [25]. We address this limitation by extending MCMAE’s backbone [25] to 3D medical images through an additional depth dimension, allowing our model to process only unmasked tokens in transformer layers. This optimization significantly improves computational efficiency and scalability, as validated in Section 4.4.1. As illustrated in Fig 5, our FPN architecture adopts MCMAE’s [25] hierarchical design, processing features at four scales (H/2×W/2 to H/16×W/16) with channel dimensions (C to 4C). Each scale utilizes StrideConv [25] downsampling and is then patchified to patches, followed by masking for unmasked patch generation. The bottom-up pathway integrates features through lateral connections and summations, producing comprehensive multi-scale representations while maintaining computational efficiency.
Figure 5: Architecture of the FPN adapted from MCMAE [25], illustrating multi-scale feature extraction and fusion through hierarchical feature maps.
Overall objective function. Our MiM introduces a hierarchically designed approach to 3D medical image representation learning through multi-level reconstruction $\mathcal{L}_{R}$ and cross-level alignment $\mathcal{L}_{C}$. We empirically set the level of masked volume to $L=3$ since the three-level masked volume can provide a good balance between the representation learning and computational efficiency. Ablation study of $L$ is on the Section 4.4.1. Specifically, the multi-level reconstruction loss in Eq. 3 can be further expanded as follows:
@@
Then, since cross-level alignment loss applied between adjacent level volumes, Eq. 6 expands as follows:
@@
where the hyper-parameters $\alpha$ are used to balance the relative contributions of these two kinds of losses. We empirically set the $\alpha$ to 0.1 based on our experiment results. The ablation study of hyperparameter $\alpha$ presents in Section 4.4.2.
@@ 4 Experiments
This section will commence by introducing the datasets and evaluation metrics. Then, we will elaborate on the implementation details of our MiM. Lastly, we will present the experimental results of MiM in comparison to existing methods, along with an analysis of our proposed approach.
4.1 Datasets and Evaluation
Pre-training datasets. To conduct a fair comparison with previous works [19, 9, 8, 22, 51, 38, 32, 52, 20], we also carried out pre-training experiments on two public datasets, i.e., BTCV [53] and TCIA Covid19 [54], and combined them to form a new dataset named 1k. Additionally, to explore the scaling ability of our proposed method compared to previous state-of-the-art methods [19, 9], we collected eight publicly accessible 3D medical image datasets consisting of 10,502 CT scans to establish our pre-training datasets, which we named 10k. It is important to note that 10k is only used for exploring the scaling ability of our proposed method, while we mainly focus on the 1k dataset for fair comparison and analysis of our proposed method. Table 1 provides a summary of the sources of each collected dataset.
Table 1: The details of each dataset in our pre-training datasets.
@@ 4.1 Datasets and Evaluation
BTCV [53] Abdomen 24
@@
LUNA16 [55] Chest 843
@@
FLARE23 [15] Abdomen 4,000
@@
HNSCC [58] Head/Neck 1,071
@@
Downstream datasets. We conduct experiments on twelve public datasets, i.e, BTCV [53], MM-WHS [60], Spleen [61], Flare22 [62], Amos22 [63], MSD Task03 [61], MSD Task06 [61], MSD Task07 [61], MSD Task08 [61], MSD Task10 [61], BrasTS 21 [61], and CC-CCII [64]. These datasets include segmentation and classification tasks, with the first ten datasets used for organ segmentation, the eleventh dataset used for lesion segmentation, the twelfth dataset used for tumor segmentation, and the last dataset used for disease classification. For BTCV [53] dataset, we adhere strictly to the data splits defined in prior studies [9, 20, 11], which comprise only training and validation sets. The training split is utilized for both pre-training and fine-tuning, while the validation split is excluded from pre-training and reserved solely for evaluation. All other datasets are unseen during pre-training. Furthermore, to evaluate the cross-modality generalization ability, we transferred the model pre-trained on CT scans to the MRI dataset BrasTS 21 [61]. We adopted consistent settings as previous work [19, 9, 20].
Evaluation metrics. Following [9, 20], we utilized Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD) to evaluate the segmentation tasks. We then utilized Accuracy (ACC) and the Area Under the Curve (AUC) to evaluate the disease classification tasks.
@@
MiM Level-1 $H^{1}W^{1}D^{1}$ $384,384,192$
MiM Level-2 $H^{2}W^{2}D^{2}$ $96,96,96$
MiM Level-3 $H^{3}W^{3}D^{2}$ $16,16,16$
MiM Resize after crop $hwd$ $96,96,96$
@@
Method Pre-traning dataset Network Dice Score(%) Avg
@@
UNETR [65] - - 93.02 94.13 94.12 66.99 70.87 96.11 77.27 89.22 82.10 70.16 76.65 65.32 59.21 79.82
Swin-UNETR [21] - - 94.06 93.54 93.80 65.51 74.60 97.09 75.94 91.80 82.36 73.63 75.19 68.00 61.11 80.53
General SSL
SimCLR [29] 1k UNETR 92.79 93.04 91.41 49.65 50.99 98.49 77.92 85.56 80.58 64.37 67.16 59.04 48.99 73.85
MoCov3 [66] 1k UNETR 91.96 92.85 92.42 68.25 72.77 94.91 78.82 88.21 81.59 71.15 75.76 66.48 58.81 79.54
@@
SimMIM [39] 1k Swin-UNETR 95.51 93.61 93.49 67.91 73.50 96.46 81.15 89.78 84.86 72.45 75.70 66.89 64.46 81.41
@@
GL-MAE [20] 1k UNETR 94.54 94.39 94.37 73.19 74.93 96.51 83.49 89.74 83.11 70.80 75.71 69.39 63.12 82.33
MAE3D [19] 1k UNETR 95.81 94.38 94.48 69.96 76.85 96.69 80.44 90.33 84.33 73.65 80.11 68.65 64.44 82.40
Rubik++ [36] 1k Swin-UNETR 96.21 90.41 89.33 75.22 72.64 97.44 79.25 89.65 83.76 74.74 78.35 67.14 61.97 81.38
SwinMM [22] 1k Swin-UNETR 94.33 94.18 94.16 72.97 74.75 96.37 83.23 89.56 82.91 70.65 75.52 69.17 62.90 81.81
@@
Figure 6: Visualization of segmentation results on BTCV validation dataset [53]. We compared MiM with Unimiss [51], PCRLv2 [52], GVSL [8], Swin-UNETR [9] and MAE3D [19].
4.2 Implementation Details
During pre-training, we followed the settings of previous works [19, 9] and provided the details of our MiM pre-training settings in Table 2. Specifically, the level-1 volume was randomly cropped from the entire CT volumes. We used the backbone of  [25] as the encoder for efficient token processing. As previous SSL [49, 20], the prediction head and projection head are implemented with MLP layers for aligning the dimension. During fine-tuning, we used Swin-UNETR [21] for fine-tuning on segmentation tasks and Swin-ViT for fine-tuning on classification tasks, as per the settings of previous works [9, 20]. Specifically, for segmentation task, we discarded the decoder and only used the backbone in the finetuning period. For classification tasks,we strictly adhered to the methodology used in prior work in general computer vision [49, 17] and 3D medical imaging [20], utilizing only the features from the final layer for pre-training. While incorporating multiscale features could enhance classification performance [45, 67, 68], we opted not to follow this approach to ensure a fair comparison with other methods. Therefore, we use only the final layer’s features, combined with a global average pooling (GAP) layer and a simple MLP classifier, to predict the category. We initialized the encoder of the network with learned parameters in the pre-training process and fine-tuned the overall network. For inference on these datasets, we applied a sliding window inference with overlapping to enable a fair comparison with previous works [9]. It’s important to note that for evaluating the pure effectiveness of our proposed method, we did not use any foundation model or post-processing techniques [7, 69].
Comparison methods. We compared our MiM method with both General SSL and Medical SSL methods. First, we compare the typical SSL methods MoCov3 [66] and MAE [19, 17], since they represent the two mainstream SSL paradigms. We also report the results of SimCLR [29] according to  [19, 20]. We further evaluate the performance of SimMIM [39], HPM [22], localMIM [19] and MCMAE [25], since they are related to our advanced hybrid MAE. We also compare with Adam [47] due to related to our hierarchical design of multi-granularity. Additionally, we compared with most existing SOTA medical SSL methods. Following common practices in SSL for natural images [70, 29, 17, 66, 49] and 3D medical images [9, 19, 20, 32, 52], we performed a single round of pre-training and fine-tuning for all methods to obtain results.